The Rule of 70 applies in any growth rate application. Let’s say you have $1000 in savings and you have three alternatives for investing these funds. How long would it take to double your savings in each of these 3 accounts? Show your answers.

• A savings account earning 1.5 % interest per year

• A U.S. Treasury bond mutual fund earning 3.5% interest per year

• A stock market mutual fund earning 9% interest per year
Answer #1

As per the rule of 70, the savings would double in:-

a) 70/1.5= 46.67 or 47 years

b) 70/3.5 = 20 years

c) 70/9 = 7.78 or 8 years